OpenStudy (anonymous):

The population of a town decreased from 15,925 people to 8759 people. What was the approximate percent decrease?

OpenStudy (anonymous):

\[\frac{15925-8759}{15925} \times 100 = \frac{7166}{15925} \times 100= 0.4499 \times 100 = 44.99 Percent\]